Barbara Muller Services (BMS) pays its employees monthly. The payroll information listed below is for January 2021, the first month of BMS's fiscal year. Assume none of the employees' earnings reached $7,000 during the month. Salaries $ 70,000 Federal income taxes to be withheld 17,000 Federal unemployment tax rate (FUTA) 0.80 % State unemployment tax rate (after FUTA deduction) 5.40 % Social security tax rate 6.2 % Medicare tax rate 1.45 % The journal entry to record payroll for the January 2021 pay period will include a debit to payroll tax expense of: Multiple Choice $5,355. $4,340. $9,695. $47,645.
Answer #1
Social security tax 4340 =70000*6.2%
Medicare tax 1015 =70000*1.45%
Federal unemployment tax 560 =70000*0.8%
State unemployment tax 3780 =70000*5.4%
Payroll tax expense 9695
$9,695 is correct option